if a recipe calls for 4oz liquid portion and we have a population of 1500 how many pounds of product would it take to feed our population?
Answer by stanbon(75887)

if a recipe calls for 4oz liquid portion per peron and we have a population of 1500 how many pounds of product would it take to feed our population?
====
Ans: 1500*4 = 6000 ounces = 6000/16 = 3750 pounds
=========================================
Cheers,
Stan H.